# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v4

@@ I-A Overview
These days, the importance of wireless sensing has grown significantly, driven by advancements in technology and the increasing demand for efficient monitoring solutions. Wireless sensing has found applications in various fields such as fall detection [2], person localization [3], and people identification [4]. Among the various manners, Wi-Fi sensing—particularly in the sub-7GHz band—stands out as a popular method with numerous advantages, as outlined in Table I. [^1] Although Wi-Fi sensing can also be implemented in the millimeter-wave band, this work focuses on sub-7GHz Wi-Fi, which is more widely deployed and better supported by existing open-source CSI extraction tools. Given that millimeter-wave Wi-Fi communication is not yet mainstream [5], all subsequent references to Wi-Fi sensing in this paper refer specifically to sub-7GHz Wi-Fi. First, Wi-Fi sensing is contactless as the object to be detected does not need to carry a device like a smartphone or smartwatch. Second, Wi-Fi sensing is privacy-preserving compared with camera. Third, Wi-Fi sensing is insensitive to occlusions and illumination conditions, as the sub-7 GHz electromagnetic waves used in Wi-Fi communication can effectively penetrate walls. Moreover, Wi-Fi sensing is ubiquitous, with over $19.5$ billion Wi-Fi devices deployed globally. With an appropriate algorithm that enables its sensing function, ubiquitous sensing function can be realized in various environments, such as homes, offices, and public spaces. Last but not least, Wi-Fi sensing has lower cost compared to dedicated radars.
@@
Functionality Accuracy Coverage Privacy Cost
Camera Comprehensive Excellent Limited (affected by occlusion/lighting) Exposed Relatively Inexpensive
Millimeter Wave Radar Fairly Comprehensive Quite Good Moderate Fairly Secure Expensive
Infrared Sensor Limited Excellent Moderate Secure Inexpensive
Sub-7 GHz Wi-Fi Fairly Comprehensive Quite Good Good (non-line-of-sight, all-day) Secure Negligible
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [6], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. Using this CSI data collection tool, a series of research employs commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [7] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purposes without sensing-specific optimization—such as low-sidelobe ambiguity functions, high time–frequency resolution, or consistent pilot structures—direct range–Doppler modeling from CSI measurements is less reliable than with radar-optimized waveforms, making it challenging to apply physical models to infer the ambient environment.
@@ I-B Cross Domain Challenge
Recently, most Wi-Fi sensing methods rely on data-driven approach via machine learning, especially for fine-grained tasks like gesture recognition [8]. However, even minor environmental changes can drastically alter the CSI patterns and significantly degrade the performance of the trained model. This environmental variability leads to a domain shift - a discrepancy between the training data distribution and real-world deployment conditions. Given the strong dependence of the training data, such domain shifts can severely undermine the generalization capability. This scenario is known as a cross-domain problem in machine learning, where the source domain (training data) and target domain (testing data) have different underlying distributions or feature spaces.
In fields like Computer Vision (CV) and Natural Language Processing (NLP), where large datasets are available in these common modalities, it is relatively easy to solve such cross-domain problems by training large models with vast amount of data [9]. However, in Wi-Fi sensing, the limited public datasets often have different formats, making it challenging to use them together to train a model. Additionally, collecting large CSI datasets on one’s own is not easy. Therefore, it is needed to develop a powerful yet adaptable small model that can quickly adapt to new environments, which has emerged as an important research direction in Wi-Fi sensing.
@@ I-C Existing Cross-Domain Wi-Fi Sensing Methods
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment Learning (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [10, 11, 12, 13, 14, 15, 16]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain.
Additionally, several cross-domain Wi-Fi sensing methods focus on few-shot learning (feature learning) [17]. These methods typically follow a paradigm where a feature extractor is first trained, and then similarities of the extracted features from the source and target domains are compared for classification. Most of these methods concentrate on efficiently designing and improving the feature extractor and similarity calculation components. For instance, Ding et al. [18] propose using a CNN trained in the source domain as a feature extractor, employing cosine similarity in the feature embedding for similarity calculations. They also implement a low-pass Butterworth filter for denoising. Building on this, AutoFi [19] is introduced (without the low-pass denoising), proposing a pre-training method to enhance the feature extractor’s capacity within the neural network. In Yang et al. [20], a different similarity calculation method is employed, specifically a Siamese network [21]. They utilize a CNN combined with Bi-LSTM as the feature extractor. Expanding upon this, CrossFi [22] uses a ResNet-based Siamese architecture, introducing a template generation method to enhance feature representation and an attention-based similarity calculation method to better capture relationships between samples. In CrossFi, the application of the Siamese network is extended to a zero-shot scenario. However, a widely recognized issue in few-shot learning methods is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [22] and Section V-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
@@ II-B Domain Shift Issue in Wi-Fi Sensing
Figure 2: UMAP Dimensionality Reduction of the WiGesture Dataset [4]: In the figure, different shapes represent different person IDs, while different colors represent different action IDs. The x and y axes correspond to the two main UMAP components.
@@
In cross-domain Wi-Fi sensing, the most significant challenge is the substantial variation in data distribution across different domains, which causes models trained in the source domain to often fail in the target domain. For better illustration, we use WiGesture dataset [4], which contains the CSI collected from 8 individuals performing 6 different gesture actions, to illustrate the domain shift problem in detail. For the gesture recognition or action classification task, we take different people as different domains. And for the people identification task, we take different actions as different domains. Fig. 2 depicts the Uniform Manifold Approximation and Projection (UMAP) based data reduction [23] (this method will be introduced in Section IV-D1). It is evident that even for the same action, individuals exhibit distinct data distributions; similarly, for the same individual, different actions show varied data distributions. These significant domain gaps can pose challenges for classification models. To illustrate this more clearly, Figs. 3a and 3b display the UMAP results for single and multiple individuals. It is apparent that within a single individual, different actions exhibit significant dissimilarities. However, when considering different individuals, the boundaries separating different actions become less distinct. These findings imply that it is challenging to generalize the learned features from one set of individuals to others. In Fig. 3c, a ResNet18 [24] is trained by setting the people with IDs 1-7 as the source domain and people with ID 0 as the target domain. It can be observed that the trained ResNet18 can distinguish different actions in the source domain but fail to distinguish different actions in the target domain. This is mainly because even for the same action, different people have distinct behavioral patterns, which can lead to significant gaps in the CSI patterns between individuals. As a result, to make the model perform well in the target domain, it is necessary to address the domain shift (cross-domain) problem.
@@
Figure 3: Data Dimension Reduction Results of the WiGesture Dataset [4]: Different colors represent different action categories. The data dimensions have been reduced to three for visualization, corresponding to the three axes in the figures. They remain consistent across subsequent figures.
@@ III Preliminary
Metric-based Method Learning-based Method Domain Alignment Method Ours Representative Methods KNN[25], K-means[26] Siamese[21], Triplet Network[27] MMD[28], GFK[29, 30] KNN-MMD Sensitivity to Quality of Support Set High Moderate None Low Stability Low Low Low High Assumption $P_{t}(y|x)=P_{s}(y|x)$ No Some methods require it[31] Yes No
As the most prevalent method in cross-domain tasks, DA encompasses various classification approaches. We primarily follow [32, 33], categorizing DA into three types: metric-based methods, learning-based methods, and DAL methods. In practice, a method may belong to multiple categories, however, this paper focuses on their primary classifications: metric-based methods refer mainly to traditional machine learning approaches that do not involve training, learning-based methods pertain to those associated with novel network architectures, and DAL methods primarily address the alignment of feature spaces between source and target domains.
@@
With the development of deep learning, there have been many learning-based methods for cross-domain tasks in recent years, including transfer learning [34], meta learning [35], FSL [36]. However, an important practical problem of these methods is that we do not know when to stop the model training. Since there is often a significant gap between the source and target domains, we cannot reliably infer the model’s performance on the target domain based on its performance on the source domain. As a result, the typical early stop strategies cannot be easily applied in such scenarios.
DAL methods aim to map the source domain and target domain to a common feature distribution, allowing the data in both domains to share the same distribution in the new space. However, most DAL methods, such as MMD [28], Kernel Mean Matching (KMM) [37], and Subspace Geodesic Flow (SGF) [38], rely on the assumption that the conditional probabilities in the source and target domains are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$, where $P_{s}$ and $P_{t}$ represent the probability distributions in the source and target domains, respectively, and $x$ and $y$ represent the data sample and ground truth category label. When training the network in the source domain, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)\approx P(y|x)$ under the source domain $P_{s}(x)$. However, since $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain $P_{t}(x)$. Consequently, these traditional methods can be viewed as attempting to find a latent space where the input data from the source and target domains are mapped to the same distribution, i.e., $P_{t}(\text{g}(x))\approx P_{s}(\text{g}(x))$, where $\text{g}(\cdot)$ is a mapping function. This allows the model to learn the parameters $\theta$ more effectively and accurately in the target domain. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not satisfied. This has been demonstrated in our experiments (Section V-C) as well as in our previous research [22], where traditional DAL methods have even failed in cross-domain Wi-Fi sensing tasks. A more detailed derivation of the formulas can be found in Section IV-B.
@@
In addition, some other researches have also identified the limitations of traditional DAL methods. Tian et al. [39] utilize KNN [25] to identify the nearest samples in the entire data domain to each sample in the source domain, expanding the distance between neighboring samples from different categories in the source domain, and shrinking the distance between neighboring samples from the target domain under the assumption that such sample pairs share the same category. However, this approach may also lead to potential failures, as shown in Fig. 1. Though the source domain samples have a relatively small distance to their neighboring samples in the target domain and a large distance to other category samples in the source domain, it cannot guarantee accurate classification in the target domain. Li et al.[40] first generate pseudo-labels for samples in the target domain using a classifier trained in the source domain and then attempt to expand the distance between samples with different categories and shrink the distance between samples with the same category. However, due to the inherent gap between the source and target domains, the pseudo-labels generated in the first step may be inaccurate, which can adversely impact the subsequent steps.
@@
Although some previous work has highlighted the disadvantages of conventional global alignment and proposed similar pseudo-label-based DAL methods, we want to emphasize the differences between our work and theirs. First, we argue that some prior DAL methods still struggle in challenging scenarios. For instance, the method proposed by [39] cannot effectively address the scenario illustrated in Fig. 1, as mentioned in the previous section. Our proposed method, which directly aligns the source and target domains within each category, represents a simple yet effective approach to achieve local alignment. The success of our method primarily relies on the quality of pseudo labels for the target domain. An inevitable issue in pseudo-label-based methods is the occurrence of incorrect classifications and noise in the data. Most previous work [40, 41, 42] first trains a network in the source domain and then uses it to classify the target domain, generating pseudo labels. Typically, these methods retain only the top classification samples from the target domain for subsequent DAL training, aiming to minimize noise and incorrect primary classifications as much as possible. While these methods have shown considerable success in fields like CV and electrocardiogram (ECG) analysis, the Wi-Fi CSI serves as a different modality, often exhibiting a completely different data distribution between source and target domains. This discrepancy can render the network trained on the source domain ineffective in the target domain. For example, using the WiGesture dataset, we train a ResNet18 in the source domain and test its performance in the target domain. The confusion matrix is shown in Fig. 4a. For category 2-4, most classification results are incorrect. Notably, in category 2, there are no correct generated labels, which hinders subsequent operations and leads to significantly low model performance. In Fig. 4b, we also display the accuracy across different categories under the top $p\%$ classification confidence level. For classes other than 0 and 4, regardless of how $p\%$ is selected, the accuracy remains low, resulting in poor quality of pseudo labels. In contrast, our method utilizes model-based KNN for pseudo label generation, which is less affected by cross-domain issues compared to neural networks. Additionally, we propose a confidence-level-based approach to mitigate the influence of noise and incorrect classifications, which has been demonstrated to be effective through sensitivity analysis. Furthermore, our method supports an early stopping mechanism, which is not feasible for most DAL methods. This feature allows us to efficiently determine when to stop the model at the optimal point, significantly enhancing model stability.

### v5

@@ I-A Overview
These days, the importance of wireless sensing has grown significantly, driven by advancements in technology and the increasing demand for efficient monitoring solutions. Wireless sensing has found applications in various fields such as fall detection [fall], person localization [zhao2024lofi], and people identification [CSI-BERT]. Among the various manners, Wi-Fi sensing—particularly in the sub-7GHz band—stands out as a popular method with numerous advantages, as outlined in Table I [zhu2025review]. [^1] Although Wi-Fi sensing can also be implemented in the millimeter-wave band, this work focuses on sub-7GHz Wi-Fi, which is more widely deployed and better supported by existing open-source CSI extraction tools. Given that millimeter-wave Wi-Fi communication is not yet mainstream [liu2024wi], all subsequent references to Wi-Fi sensing in this paper refer specifically to sub-7GHz Wi-Fi. First, Wi-Fi sensing is contactless as the object to be detected does not need to carry a device like a smartphone or smartwatch. Second, Wi-Fi sensing preserves privacy by not capturing visual images or personally identifiable information, in contrast to cameras. Moreover, the raw data it collects is not directly interpretable by humans, further protecting privacy. Third, Wi-Fi sensing is insensitive to occlusions and illumination conditions, as the sub-7 GHz electromagnetic waves used in Wi-Fi communication can effectively penetrate walls. Moreover, Wi-Fi sensing is ubiquitous, with over $19.5$ billion Wi-Fi devices deployed globally. With an appropriate algorithm that enables its sensing function, ubiquitous sensing function can be realized in various environments, such as homes, offices, and public spaces. Last but not least, Wi-Fi sensing has lower cost compared to dedicated radars.
@@
Functionality Accuracy Coverage Privacy Cost Camera Highly versatile High Limited (affected by occlusion/lighting) Low Low to moderate Millimeter Wave Radar Moderately versatile High Moderate High Expensive Infrared Sensor Limited Medium Moderate High Low Sub-7 GHz Wi-Fi Moderately versatile Medium Good (non-line-of-sight, all-day) High Very low
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [halperin2011tool], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. Using this CSI data collection tool, a series of research employs commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [802.11bf] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purposes without sensing-specific optimization—such as low-sidelobe ambiguity functions, high time–frequency resolution, or consistent pilot structures—direct range–Doppler modeling from CSI measurements is less reliable than with radar-optimized waveforms, making it challenging to apply physical models to infer the ambient environment.
@@ I-B Cross Domain Challenge
Recently, most Wi-Fi sensing methods rely on data-driven approach via machine learning, especially for fine-grained tasks like gesture recognition [nirmal2021deep]. However, even minor environmental changes can drastically alter the CSI patterns and significantly degrade the performance of the trained model. This environmental variability leads to a domain shift - a discrepancy between the training data distribution and real-world deployment conditions. Given the strong dependence of the training data, such domain shifts can severely undermine the generalization capability. This scenario is known as a cross-domain problem in machine learning, where the source domain (training data) and target domain (testing data) have different underlying distributions or feature spaces.
In fields like Computer Vision (CV) and Natural Language Processing (NLP), where large datasets are available in these common modalities, it is relatively easy to solve such cross-domain problems by training large models with vast amount of data [chen2024overview]. However, in Wi-Fi sensing, the limited public datasets often have different formats, making it challenging to use them together to train a model. Additionally, collecting large CSI datasets on one’s own is not easy. Therefore, it is needed to develop a powerful yet adaptable small model that can quickly adapt to new environments, which has emerged as an important research direction in Wi-Fi sensing.
@@ I-C Existing Cross-Domain Wi-Fi Sensing Methods
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment Learning (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [MMD1, MMD2, MMD3, MMD4, MMD5, SN-MMD, airfi]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain.
Additionally, several cross-domain Wi-Fi sensing methods focus on few-shot learning (feature learning) [wang2025survey]. These methods typically follow a paradigm where a feature extractor is first trained, and then similarities of the extracted features from the source and target domains are compared for classification. Most of these methods concentrate on efficiently designing and improving the feature extractor and similarity calculation components. For instance, Ding et al. [ding2021wi] propose using a CNN trained in the source domain as a feature extractor, employing cosine similarity in the feature embedding for similarity calculations. They also implement a low-pass Butterworth filter for denoising. Building on this, AutoFi [AutoFi] is introduced (without the low-pass denoising), proposing a pre-training method to enhance the feature extractor’s capacity within the neural network. In Yang et al. [yang2019learning], a different similarity calculation method is employed, specifically a Siamese network [siamese]. They utilize a CNN combined with Bi-LSTM as the feature extractor. Expanding upon this, CrossFi [CSi-Net] uses a ResNet-based Siamese architecture, introducing a template generation method to enhance feature representation and an attention-based similarity calculation method to better capture relationships between samples. In CrossFi, the application of the Siamese network is extended to a zero-shot scenario. However, a widely recognized issue in few-shot learning methods is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [CSi-Net] and Section V-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
@@ II-B Domain Shift Issue in Wi-Fi Sensing
Figure 2: UMAP Dimensionality Reduction of the WiGesture Dataset [CSI-BERT]: In the figure, different shapes represent different person IDs, while different colors represent different action IDs. The x and y axes correspond to the two main UMAP components.
@@
In cross-domain Wi-Fi sensing, the most significant challenge is the substantial variation in data distribution across different domains, which causes models trained in the source domain to often fail in the target domain. For better illustration, we use WiGesture dataset [CSI-BERT], which contains the CSI collected from 8 individuals performing 6 different gesture actions, to illustrate the domain shift problem in detail. For the gesture recognition or action classification task, we take different people as different domains. And for the people identification task, we take different actions as different domains. Fig. 2 depicts the Uniform Manifold Approximation and Projection (UMAP) based data reduction [UMAP] (this method will be introduced in Section IV-D1). It is evident that even for the same action, individuals exhibit distinct data distributions; similarly, for the same individual, different actions show varied data distributions. These significant domain gaps can pose challenges for classification models. To illustrate this more clearly, Figs. 3a and 3b display the UMAP results for single and multiple individuals. It is apparent that within a single individual, different actions exhibit significant dissimilarities. However, when considering different individuals, the boundaries separating different actions become less distinct. These findings imply that it is challenging to generalize the learned features from one set of individuals to others. In Fig. 3c, a ResNet18 [ResNet] is trained by setting the people with IDs 1-7 as the source domain and people with ID 0 as the target domain. It can be observed that the trained ResNet18 can distinguish different actions in the source domain but fail to distinguish different actions in the target domain. This is mainly because even for the same action, different people have distinct behavioral patterns, which can lead to significant gaps in the CSI patterns between individuals. As a result, to make the model perform well in the target domain, it is necessary to address the domain shift (cross-domain) problem.
@@
Figure 3: Data Dimension Reduction Results of the WiGesture Dataset [CSI-BERT]: Different colors represent different action categories. The data dimensions have been reduced to three for visualization, corresponding to the three axes in the figures. They remain consistent across subsequent figures.
@@ III Preliminary
Metric-based Method Learning-based Method Domain Alignment Method Ours Representative Methods KNN[KNN], K-means[kmeans] Siamese[siamese], Triplet Network[triplet] MMD[MMD], GFK[GFK, zhang2018grassmannian] KNN-MMD Sensitivity to Quality of Support Set High Moderate None Low Stability Low Low Low High Assumption $P_{t}(y|x)=P_{s}(y|x)$ No Some methods require it[DANN] Yes No
As the most prevalent method in cross-domain tasks, DA encompasses various classification approaches. We primarily follow [zhu2024cross, he2024domain], categorizing DA into three types: metric-based methods, learning-based methods, and DAL methods. In practice, a method may belong to multiple categories, however, this paper focuses on their primary classifications: metric-based methods refer mainly to traditional machine learning approaches that do not involve training, learning-based methods pertain to those associated with novel network architectures, and DAL methods primarily address the alignment of feature spaces between source and target domains.
@@
With the development of deep learning, there have been many learning-based methods for cross-domain tasks in recent years, including transfer learning [transfer], meta learning [meta], FSL [few_shot]. However, an important practical problem of these methods is that we do not know when to stop the model training. Since there is often a significant gap between the source and target domains, we cannot reliably infer the model’s performance on the target domain based on its performance on the source domain. As a result, the typical early stop strategies cannot be easily applied in such scenarios.
DAL methods aim to map the source domain and target domain to a common feature distribution, allowing the data in both domains to share the same distribution in the new space. However, most DAL methods, such as MMD [MMD], Kernel Mean Matching (KMM) [KMM], and Subspace Geodesic Flow (SGF) [SGF], rely on the assumption that the conditional probabilities in the source and target domains are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$, where $P_{s}$ and $P_{t}$ represent the probability distributions in the source and target domains, respectively, and $x$ and $y$ represent the data sample and ground truth category label. When training the network in the source domain, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)\approx P(y|x)$ under the source domain $P_{s}(x)$. However, since $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain $P_{t}(x)$. Consequently, these traditional methods can be viewed as attempting to find a latent space where the input data from the source and target domains are mapped to the same distribution, i.e., $P_{t}(\text{g}(x))\approx P_{s}(\text{g}(x))$, where $\text{g}(\cdot)$ is a mapping function. This allows the model to learn the parameters $\theta$ more effectively and accurately in the target domain. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not satisfied. This has been demonstrated in our experiments (Section V-C) as well as in our previous research [CSi-Net], where traditional DAL methods have even failed in cross-domain Wi-Fi sensing tasks. A more detailed derivation of the formulas can be found in Section IV-B.
@@
In addition, some other researches have also identified the limitations of traditional DAL methods. Tian et al. [KNN-MMD] utilize KNN [KNN] to identify the nearest samples in the entire data domain to each sample in the source domain, expanding the distance between neighboring samples from different categories in the source domain, and shrinking the distance between neighboring samples from the target domain under the assumption that such sample pairs share the same category. However, this approach may also lead to potential failures, as shown in Fig. 1. Though the source domain samples have a relatively small distance to their neighboring samples in the target domain and a large distance to other category samples in the source domain, it cannot guarantee accurate classification in the target domain. Li et al.[EEG] first generate pseudo-labels for samples in the target domain using a classifier trained in the source domain and then attempt to expand the distance between samples with different categories and shrink the distance between samples with the same category. However, due to the inherent gap between the source and target domains, the pseudo-labels generated in the first step may be inaccurate, which can adversely impact the subsequent steps.
@@
Although some previous work has highlighted the disadvantages of conventional global alignment and proposed similar pseudo-label-based DAL methods, we want to emphasize the differences between our work and theirs. First, we argue that some prior DAL methods still struggle in challenging scenarios. For instance, the method proposed by [KNN-MMD] cannot effectively address the scenario illustrated in Fig. 1, as mentioned in the previous section. Our proposed method, which directly aligns the source and target domains within each category, represents a simple yet effective approach to achieve local alignment. The success of our method primarily relies on the quality of pseudo labels for the target domain. An inevitable issue in pseudo-label-based methods is the occurrence of incorrect classifications and noise in the data. Most previous work [EEG, ge2023unsupervised, zhu2020deep] first trains a network in the source domain and then uses it to classify the target domain, generating pseudo labels. Typically, these methods retain only the top classification samples from the target domain for subsequent DAL training, aiming to minimize noise and incorrect primary classifications as much as possible. While these methods have shown considerable success in fields like CV and electrocardiogram (ECG) analysis, the Wi-Fi CSI serves as a different modality, often exhibiting a completely different data distribution between source and target domains. This discrepancy can render the network trained on the source domain ineffective in the target domain. For example, using the WiGesture dataset, we train a ResNet18 in the source domain and test its performance in the target domain. The confusion matrix is shown in Fig. 4a. For category 2-4, most classification results are incorrect. Notably, in category 2, there are no correct generated labels, which hinders subsequent operations and leads to significantly low model performance. In Fig. 4b, we also display the accuracy across different categories under the top $p\%$ classification confidence level. For classes other than 0 and 4, regardless of how $p\%$ is selected, the accuracy remains low, resulting in poor quality of pseudo labels. In contrast, our method utilizes model-based KNN for pseudo label generation, which is less affected by cross-domain issues compared to neural networks. Additionally, we propose a confidence-level-based approach to mitigate the influence of noise and incorrect classifications, which has been demonstrated to be effective through sensitivity analysis. Furthermore, our method supports an early stopping mechanism, which is not feasible for most DAL methods. This feature allows us to efficiently determine when to stop the model at the optimal point, significantly enhancing model stability.